### v1

# BESSTIE: A Benchmark for Sentiment and Sarcasm Classification for Varieties of English

Abstract.
Despite large language models (LLMs) being known to exhibit bias against non-mainstream varieties, there are no known labeled datasets for sentiment analysis of English. To address this gap, we introduce BESSTIE, a benchmark for sentiment and sarcasm classification for three varieties of English: Australian (en-AU), Indian (en-IN), and British (en-UK). Using web-based content from two domains, namely, Google Place reviews and Reddit comments, we collect datasets for these language varieties using two methods: location-based and topic-based filtering. Native speakers of the language varieties manually annotate the datasets with sentiment and sarcasm labels. Subsequently, we fine-tune nine large language models (LLMs) (representing a range of encoder/decoder and mono/multilingual models) on these datasets, and evaluate their performance on the two tasks. Our results reveal that the models consistently perform better on inner-circle varieties (i.e., en-AU and en-UK), with significant performance drops for en-IN, particularly in sarcasm detection. We also report challenges in cross-variety generalisation, highlighting the need for language variety-specific datasets such as ours. BESSTIE promises to be a useful evaluative benchmark for future research in equitable LLMs, specifically in terms of language varieties. The BESSTIE datasets, code, and models are currently available on request, while the paper is under review. Please email aditya.joshi@unsw.edu.au.
bias, language variety, dialectal bias, language models, sentiment classification, sarcasm classification, benchmark, natural language processing
ccs: Computing methodologies Natural language processingccs: Computing methodologies Language resources
1. Introduction
Benchmark-based evaluation (Socher et al., 2013) of large language models (LLMs) is the prevalent norm in natural language processing (NLP). Benchmarks provide labeled datasets (Wang et al., 2019; Muennighoff et al., 2023) for specific tasks so that LLMs can be evaluated against them. However, most NLP benchmarks do not contain text in language varieties such as national varieties, dialects, sociolects or creoles (Plank, 2022; Lent et al., 2024; Joshi et al., 2024). As a result, despite the superlative performance of LLMs on Standard American English, LLMs are not evaluated on other varieties of English. Knowing that LLMs can be biased against certain language varieties (as shown by  (Deas et al., 2023) for African-American English), little empirical evidence is found for their bias towards several language varieties, particularly those of English. An exception is Multi-VALUE (Ziems et al., 2023) which creates synthetic datasets of language varieties using linguistically-informed syntactic transformations and reports a degraded performance on the GLUE benchmark. However, such synthetic texts are insufficient in measuring language variety bias because language varieties are more than syntax and encompass orthography, vocabulary, and cultural pragmatics (Nguyen, 2021). As a result, labeled datasets of natural text are crucial to measure the bias in LLMs towards these varieties.
Benchmark Sent. Sarc. Eng. Var.
Cieliebak et al. (2017) (Cieliebak et al., 2017)
Wang et al. (2018) (Wang et al., 2018)
Alharbi et al. (2020) (Alharbi et al., 2020)
Abu Farha et al. (2021) (Abu Farha et al., 2021)
Elmadany et al. (2023) (Elmadany et al., 2023)
Faisal et al. (2024) (Faisal et al., 2024)
[…]
Table 1. Comparison of BESSTIE with past benchmarks for sentiment or sarcasm classification. ‘Sent.’ indicates sentiment classification, ‘Sarc.’ denotes sarcasm classification, ‘Eng.’ denotes English, and ‘Var.’ denotes language varieties. A checkmark (✓) denotes the availability of a particular feature, while a cross (✕) indicates its absence.
To address this gap, we introduce BESSTIE, A BEnchmark for Sentiment and Sarcasm classification for varieTIes of English, namely, Australian (en-AU), Indian (en-IN), and British (en-UK). BESSTIE comprises a manually labeled dataset of textual posts along with performance results of fine-tuned LLMs on the two classification tasks. The dataset consists of text samples collected from two web-based domains, namely, Google Places reviews, and Reddit comments, using two filtering methods: location-based and topic-based filtering respectively. Following that, we obtain manual annotations for boolean sentiment and sarcasm labels from native speakers of the language varieties. The novelty of the BESSTIE dataset can be seen in Table 1. BESSTIE improves not only upon popular sentiment/sarcasm-labeled datasets but also a recent dialectal benchmark, DialectBench (Faisal et al., 2024), which contains dialectal datasets for sentiment classification but for languages other than English. Therefore, the BESSTIE dataset is novel in the introduction of new language varieties (specifically, varieties of English) and, an additional task, sarcasm classification.
We then use the dataset to model both the tasks (sentiment and sarcasm classification) as binary classification problems. We evaluate nine fine-tuned LLMs, spanning encoder/decoder and monolingual/multilingual models. Our results indicate that current models exhibit a significant decrease in performance when processing text written in en-IN, an outer-circle variety, in comparison with inner-circle varieties of English [^1] Inner-circle English varieties are those spoken by people who use English as a first language. Outer-circle varieties are those spoken by bi/multilingual speakers. (en-AU and en-UK). The performance on sarcasm classification is consistently low across the three varieties, providing pointers for future work. Following that, we examine the aggregate performance for the three varieties, the two tasks and model attributes (encoder/decoder and monolingual/multilingual). We also examine cross-variety and cross-domain performance on all models highlighting the need for BESSTIE. BESSTIE evaluation provides a solid baseline to examine biases of contemporary LLMs, and will help to develop new methods to mitigate these biases.
The contributions of this work are:
  • Creation of a manually annotated dataset of user-generated content with sentiment and sarcasm labels for three varieties of English, called BESSTIE.
  • Evaluating nine LLMs on BESSTIE and analysing their performance to identify challenges arising due to language varieties.
An overview of BESSTIE is described in Figure 1. The paper describes the dataset (Section 4; referring to ‘Collection’ and ‘Annotation’ in the figure) and the evaluation methodology (Section 5; ‘Evaluation’ in the Figure) of BESSTIE. The experiment setup and results are in Sections 6 and 7 respectively, followed by error analysis in Section 8.
2. Related work
NLP Benchmarks such as GLUE (Wang et al., 2018), SuperGLUE (Wang et al., 2019), and DynaBench (Kiela et al., 2021) have played a pivotal role in evaluating LLMs on multiple NLP tasks, including sentiment classification. However, these benchmarks are limited in their ability to capture the linguistic nuances of non-standard language varieties. Efforts to address this include datasets for African languages and dialects (Muhammad et al., 2023), Arabic dialects (Elmadany et al., 2023), and CreoleVal (Lent et al., 2023) which provides datasets for 28 creole languages. More recently, DIALECTBENCH (Faisal et al., 2024) is a collection of dialectal datasets, and associated NLP models across 281 dialectal varieties for several tasks. However, the benchmark does not contain English dialectal datasets for sentiment and sarcasm classification. BESSTIE is the first benchmark for sentiment and sarcasm classification specifically for non-standard varieties of English.
Similarly, several sarcasm classification datasets for standard variety of English exist, consisting of reddit comments (Khodak et al., 2018), amazon reviews (Filatova, 2012), and tweets (Ptáček et al., 2014; Painter et al., 2022; Abercrombie and Hovy, 2016; Oprea and Magdy, 2020). However, there is a notable absence of sarcasm classification datasets that account for non-standard varieties of English. ArSarcasm-v2 (Abu Farha et al., 2021) is a dataset annotated for sarcasm, sentiment, and the dialect of each tweet for Arabic. Abu Farha and Magdy (2021) benchmarked Arabic-specific transformer models on this dataset, noting that models trained on Arabic data, including varieties of the language, tend to perform better.
We emphasise that the papers listed above approach benchmark creation in a similar manner as BESSTIE: they create an annotated dataset for a task, and test a broad range of models to highlight the potential for future work. In that sense, these benchmark papers acknowledge that creating a dataset and measuring bias is an important task achieved by papers that introduce benchmarks, without necessarily presenting a new method itself for the task. This holds true for our paper as well.
3. Ethical Considerations
The project received ethics approval from the Human Research Ethics Committee at UNSW, Sydney (reference number: iRECS6514). We treat the reviews as separate texts and do not aggregate them by user in any way. The Reddit comments and Google reviews used in this paper are from publicly available posts, accessible through official APIs. We adhere to the terms of service and policies outlined by these platforms while gathering data. We do not attempt to de-identify any users or collect aggregate information about individual users. The dataset, including sentiment and sarcasm annotations, is shared in a manner that preserves user privacy and abides by the rules and guidelines of the source platforms.
Figure 1. Methodology of BESSTIE. (a) Collection is detailed in Section 4.1, (b) Annotation is detailed in 4.2, and (c) Evaluation is detailed in Section 5.
4. Dataset Creation
We now describe parts (a) and (b) of Figure 1 which result in a manually annotated dataset for sentiment and sarcasm classification.
4.1. Data Collection
We collect textual posts from two domains: Google Places reviews and Reddit comments, using two filtering methods: location-based and topic-based respectively. Location-based filtering implies that we select reviews that were posted for locations in the three countries. Specifically, we collect reviews posted in cities in AU, IN and UK, and their corresponding ratings ($1$ to $5$ stars, where $5$ is the highest.) using the Google Places API [^2] https://developers.google.com/maps/documentation/places/web-service/overview; Accessed on 23 July 2024 from the place types, also defined by the API. The criteria for city selection are based on population thresholds specific to each country (en-AU: $20$K; en-IN:$100$K; en-UK: $50$K). We further filter the dataset to only contain reviews with closer ratings ($2$ and $4$ stars). We also filter out any non-English content using language probabilities calculated by fastText (Grave et al., 2018) word vectors, using a threshold of $0.98$. Additionally, to mitigate the risk of including reviews written by tourists rather than residents, we exclude reviews from locations designated as ‘tourist attractions’ [^3] https://developers.google.com/maps/documentation/places/web-service/supported_types; Accessed on 13 November 2024. by the Google Places API.
To create the reddit subset, we employ topic-based filtering and choose up to four subreddits per variety (en-AU: ‘melbourne’, ‘AustralianPolitics’,‘AskAnAustralian’; en-IN: ‘India’, ‘IndiaSpeaks’, ‘BollyBlindsNGossip’; en-UK: ‘England’, ‘Britain’, ‘UnitedKingdom’, ‘GreatBritishMemes’). We select these subreddits based on the understanding that they feature popular, discussions specific to a variety, making it highly likely that contributors use the appropriate language variety. For each variety, we scrape $12,000$ comments evenly across the selected subreddits, capping at $20$ comments per post and focusing on recent posts. These are then randomly sampled and standardized to $3,000$ comments per variety before manual annotation.
4.2. Annotation
We hire three annotators [^4] Two of these are the authors of the paper who were also involved in the evaluation. each for the three language varieties. The annotators assign the processed reviews and posts two labels: sentiment and sarcasm. The choice of labels given are negative, positive, and discard. We instruct the annotators to use the discard label for uninformative examples with no apparent polarity and in a few cases, for computer-generated messages. We compensate the annotators at the casual employment rate prescribed by the host institution.
Figure 2. Confusion matrix showing the overlap in variety annotations between annotators, ant-IN, and ant-UK. Rows represent the labels assigned by ant-IN, and columns represent the labels assigned by ant-UK. The principle diagonal elements indicate agreement between annotators, while off-diagonal elements highlight disagreements.
4.3. Data Quality & Statistics
Data Quality
While previous studies utilise location-based filtering to evaluate the robustness of language technologies across varieties (Blodgett et al., 2016; Sun et al., 2023), it needs to be examined if the collected text is indeed from these national varieties. To address this, we conduct an annotation exercise with en-IN and en-UK annotators, asking them to manually identify the variety of a given text. Specifically, they label a random sample of $300$ texts ($150$ from reviews and $150$ from comments) as en-AU, en-IN, en-UK, or Cannot say. Annotator agreement is measured using Cohen’s kappa ($\kappa$).
With respect to the true label (based on the location or topic), results show higher agreement with the en-IN annotator ($\kappa$ = 0.41) compared to the en-UK annotator ($\kappa$ = 0.34), indicating fair to moderate consistency. The inter-annotator agreement itself is $0.26$. Figure 2 shows that annotators find it difficult to agree on inner-circle varieties, and have the highest agreement when identifying en-IN variety, with 46 agreements. While agreement between annotators is limited for inner-circle varieties, the higher agreement for the en-IN indicates that the subset reliably represents this variety, thereby validating our collection methodology. As a result, the two subsets (reviews and comments) collectively form a good representative sample for language varieties.
Note: We also attempt to fine-tune an encoder model fine-tuned on an unlabeled dialectal corpus, to predict the language varieties. However, this was unsuccessful due to: (a) unavailability of the dataset for en-UK, (b)the dataset being dated and inaccurate in the domain of the collected data.
Variety Comment Sentiment Sarcasm
en-AU Well set out, good stock and very friendly country staff 1 0
en-AU S**t. Christ isn’t back, is he? 0 1
en-IN Good quality foods are supplied here 1 0
en-IN Coz we all have free internet. 0 1
en-UK Traditional friendly pub. Excellent beer 1 0
en-UK What a brave potatriot 0 1
Table 2. Reviews and comments from different varieties with their corresponding annotated sentiment and sarcasm labels.
Dataset Statistics
The benchmark dataset contains annotations for two tasks: sentiment and sarcasm classification, distributed across the three varieties. Table 3 provides a breakdown of the annotated dataset (consisting of google and reddit), including the number of samples in each split (Train, Validation, and Test), along with average word length and language probability. The % Positive sentiment and % Positive sarcasm indicate the proportion of samples with positive sentiment and true sarcasm respectively. These values are the same for train, validation and test set since we perform stratified sampling. We present a few example comments along with their corresponding sentiment and sarcasm labels in Table 2. The examples illustrate the diversity of language use and the challenges involved in detecting both sentiment and sarcasm across the varieties.
Variety Subset Train Valid Test % Positive Sentiment % Positive Sarcasm $P(\bar{\texttt{eng}})$ $\bar{w}$
en-AU google 946 130 270 73% 7% 0.99 63.97
reddit 1763 241 501 32% 42% 0.98 51.72
en-IN google 1648 225 469 75% 1% 0.99 44.34
reddit 1686 230 479 25% 13% 0.87 26.92
en-UK google 1817 248 517 75% 0% 0.99 72.21
reddit 1007 138 287 12% 22% 0.98 38.04
Table 3. Dataset statistics for each variety, subset, split, and label type. $\mu$ represents the average probability of English. $P(\bar{\texttt{eng}})$ is the average language probability (calculated using fastText (Grave et al., 2018) word vectors) of a review or comment being in English. $\bar{w}$ denotes average number of words in a review/comment.
5. Evaluation Methodology
With the annotated dataset in place, we now discuss the evaluation methodology used to evaluate the LLMs on the dataset.
The training strategies for encoder and decoder models are as follows. We train encoder models using cross-entropy loss, weighted on class distribution:
$\mathcal{L}_{\texttt{{Encoder}}}=-\frac{1}{N}\sum_{j=1}^{N}\sum_{k=1}^{C}w_{k}\cdot y_{j,k}\cdot\log p_{j,k}$
Here, $N$ denotes the total number of samples in the batch, $C$ represents the number of classes (2 since the tasks are binary classification), and $w_{k}$ is the weight assigned to class k, calculated based on the class distribution ($w_{\texttt{neg}}$ for the negative and $w_{\texttt{pos}}$ for the positive class). $y_{j}$ is the expected output label for the $j$th instance where $y_{j,1}=1$ if the label is 1 and $y_{j,0}=1$ if the label is 0. $p_{j,k}$ is the probability of the $j$th instance for $k$th class.
In contrast, for decoder models, we perform zero-shot instruction fine-tuning (Ouyang et al., 2022), using maximum likelihood estimation – a standard learning objective for causal language modeling (Jain et al., 2023):
$\mathcal{L}_{\texttt{{Decoder}}}=-\frac{1}{N}\sum_{j=1}^{N}\left\{\sum_{i=|\mathbf{x}^{j}|+1}^{|\mathbf{x}^{j}|+|\mathbf{t}^{j}|}\log p(\mathbf{x}_{i}^{j}|\mathbf{x}_{<i}^{j})\right\}$
Here, $\mathbf{x}_{<i}^{j}=[\mathbf{x}_{1}^{j},\ldots,\mathbf{x}_{i-1}^{j}]$ denotes the subsequence before $\mathbf{x}_{i}^{j}$ and ||$\cdot$|| ⋅ | is the number of tokens. Specifically, $|t_{j}|$ is the number of generated tokens and is set to $1$. In other words, given the text and a task-specific prompt (given in Table 4555The forms are experimentally determined using a few test examples.), the expected behaviour of the LLM is to generate either 1 (for positive) or 0 (for negative).
Task Prompt
Sentiment Classification “Generate the sentiment of the given text. Do not give an explanation.”
Sarcasm Classification “Predict if the given text is sarcastic. Do not give an explanation.”
Table 4. Task-specific prompts used to fine-tune decoder models. Both prompts restrict the model from generating an explanation.
6. Experiment Details
We conduct our experiments on a total of nine LLMs. These include six encoder models, three trained on English corpora: BERT-Large (bert(Devlin et al., 2019), RoBERTa-Large (roberta(Liu et al., 2019), ALBERT-XXL-v2 (albert(Lan et al., 2020); and three multilingual models: Multilingual-BERT-Base (mbert), Multilingual-DistilBERT-Base (mdistil(Sanh et al., 2020), XLM-RoBERTa-Large (xlm-r(Goyal et al., 2021). We also evaluate three decoder models, one trained on English corpora: Gemma2-27B-Instruct (gemma(Gemma Team et al., 2024) and two multilingual models: Mistral-Small-Instruct-2409 (mistral(Jiang et al., 2023), Qwen2.5-72B-Instruct (qwen(Yang et al., 2024). The model-specific details including architecture, language coverage of pre-training corpus, and number of parameters are described in Table 5.
While encoder models are fine-tuned with full precision, we fine-tune quantized decoder models using QLoRA (Dettmers et al., 2023) adapters, targeting all linear layers. All models are fined-tuned for 30 epochs, with a batch size of 8 and Adam optimiser. We choose an optimal learning rate by performing a grid search over the following values: 1e-5, 2e-5, and 3e-5. All experiments are performed using two NVIDIA A100 80GB GPUs.
We report our task performances on three macro-averaged F-Score, i.e., unweighted average, disregarding the imbalance in label distribution.
Model Arch. Lang. # Params
albert Encoder English 223 M
bert Encoder English 340 M
roberta Encoder English 355 M
mbert Encoder Multilingual 177 M
mdistil Encoder Multilingual 134 M
xlm-r Encoder Multilingual 355 M
[…]
mistral Decoder Multilingual 22 B
qwen Decoder Multilingual 72 B
Table 5. Details of models including architecture (Arch.), language coverage in pretraining data (Lang.), and number of trainable parameters (# Params). Here, M is millions and B is billion.
7. Results
Figure 3. Performance comparison of various models on the sentiment classification task across different English varieties (en-AU, en-IN, and en-UK).
7.1. Overall results
We first present our results on the two tasks where models are trained and evaluated on the same variety. Figure 3 and 4 describe the model performances, reported using F-Score, on the sentiment and sarcasm classification tasks respectively. While sentiment classification is performed using both google and reddit subsets, due to the absence of sarcasm labels in google subset (as shown in Table 3), sarcasm classification is conducted only on reddit subset.
Sentiment Classification
Table 6 shows that, for sentiment classification, mistral achieves the highest average performance across all varieties, with an F-Score of 0.91 on the google subset and 0.84 on the reddit subset. In contrast, qwen reports the lowest average performance across all varieties on the google subset, with an average F-Score of 0.45, while gemma exhibits the lowest average performance across all varieties on the reddit subset, with an F-Score of 0.60.
Figure 4. Performance comparison of various models on the sarcasm classification task across different English varieties (en-AU, en-IN, and en-UK).
Sarcasm Classification
Similarly, Table 7 shows that, for sarcasm classification, both mistral and mbert achieve the highest average performance across all varieties, with an F-Score of 0.68. In contrast, gemma again reports the lowest average performance across all varieties, with an average F-Score of 0.45. The lower performance on sarcasm classification as compared to sentiment classification is expected because sarcasm often relies on not just subtle linguistic cues but also localised contemporary contextual information (Abercrombie and Hovy, 2016).
Model Enco. Mono. google-Sentiment reddit-Sentiment reddit-Sarcasm
albert 0.88 0.75 0.65
bert 0.88 0.80 0.60
roberta 0.89 0.79 0.55
mbert 0.88 0.65 0.68
mdistil 0.86 0.78 0.51
xlm-r 0.77 0.73 0.62
gemma 0.79 0.60 0.45
mistral 0.91 0.84 0.68
qwen 0.45 0.82 0.52
[…]
Table 6. Model performances, averaged across all varieties, on sentiment and sarcasm classification tasks across two subsets (google and reddit). $\mu$ represents the average performance across all models. ‘Enco.’ represents encoder models. ‘Mono.’ represents monolingual models.
7.2. Impact of Domain and Models
The average results for the models highlight the need to further probe into factors influencing model performances, namely, domain, model properties and language varieties. The results in Tables 7 and 8 describe the model performances (reporting F-Score), grouped based on model properties. Finally, task performances, averaged across all models, on varieties are presented in Table 9.
google vs. reddit
Our findings (the last row $\mu$ in Table 6), indicate that models perform better on google subset, achieving an F-Score of 0.81, compared to an F-Score of 0.75 on reddit subset. This difference can be explained by the distinct writing styles of the two sources. Reviews in google are generally more formal and informative, while posts and comments from reddit often exhibit a short (as shown in Table 3), conversational tone typical of social media. Additionally, posts and comments in reddit come from forums frequented by local speakers, which means the language and expressions used are more reflective of the variety and cultural nuances.
Domain-Task Enco. Deco.
google-Sentiment 0.86 0.72
reddit-Sentiment 0.75 0.75
reddit-Sarcasm 0.60 0.55
[…]
Table 7. Task performances, averaged across all varieties, grouped based on model architectures. $\mu$ represents the average performance across all domain-task pairs.‘Enco.’ represents encoder models while ‘Deco.’ represents decoder models.
Domain-Task Mono. Mult.
google-Sentiment 0.86 0.78
reddit-Sentiment 0.73 0.76
reddit-Sarcasm 0.56 0.60
[…]
Table 8. Task performances, averaged across all varieties, grouped based on language diversity in pre-training data. ‘Mono.’ represents monolingual models, and ‘Mult.’ represents multilingual models. $\mu$ represents the average performance across all domain-task pairs.
Model properties
Table 7 shows a consistent trend where encoder models report higher performance than decoder models across different tasks. This performance gap is expected as encoder architecture is inherently better suited for sequence classification tasks, while decoder architecture is optimised for text generation tasks. Table 8 shows that, although marginal, monolingual models report a higher average performance compared to multilingual models. This performance gap is prominent on the task involving google subset, while multilingual models perform better on tasks involving reddit subset.
Domain-Task en-AU en-IN en-UK
google-Sentiment 0.94 0.64 0.86
reddit-Sentiment 0.78 0.69 0.78
reddit-Sarcasm 0.62 0.56 0.58
[…]
Table 9. Task performances, averaged across all models, on all varieties. $\mu$ represents the average performance across all domain-task pairs.
Comparison of language varieties
Finally, Table 9 shows that models, on average, perform better on both tasks when trained on en-AU variety, closely followed by performances on en-UK variety. Both en-AU and en-UK are inner-circle varieties. Conversely, models consistently perform the worst on en-IN, an outer-circle variety.
Figure 5. Cross-variety performance analysis of mistral. The figure compares three different scenarios: pre-trained (PT), in-variety fine-tuning, and cross-variety fine-tuning for sentiment and sarcasm classification across all varieties.
Figure 6. Cross-variety performance analysis of bert. The figure compares two different scenarios: in-variety fine-tuning, and cross-variety fine-tuning for sentiment and sarcasm classification across all varieties.
Figure 7. Cross-domain performance analysis of mistral and bert comparing a decoder and an encoder model.
7.3. Cross-variety/domain evaluation
Finally, we perform cross-variety/cross-domain evaluation. This refers to the scenario where the datasets used to train and test a model are from different groups (i.e., trained on en-AU, tested on en-IN, for example). The following discussion is focused on generalisation: “can models trained on one language variety generalise well to others”.
Cross-variety evaluation
We report the cross-variety results of our best-performing model, mistral. Figure 5 compares the performance of mistral across three variants: pre-trained, in-variety fine-tuning, and cross-variety fine-tuning, for the three dataset subsets using colour-coded matrices. The rows indicate the subset that the model is trained on (i.e., en-AU, en-IN, and en-UK), along with PT, indicating that the pre-trained mistral is used. Similarly, the columns indicate the test subset (i.e., en-AU, en-IN, and en-UK). The pre-trained variant of the model (the first row in each matrix) achieves high F-Score for en-AU and en-UK, for the sentiment classification task. Fine-tuning the model improves in-variety performance (shown by the principal diagonal of the matrices), with noticeable gains in F-Score on both subsets. The model performance on cross-variety evaluation remains relatively stable across different varieties of English, indicating that domain-specific data plays a minor role in influencing cross-variety generalisation for sentiment classification.
As seen earlier, the model performs significantly worse on sarcasm classification compared to sentiment classification. While in-variety evaluation shows that fine-tuning improves over the pre-trained model with substantial increases in F-Score, it negatively impacts cross-variety generalization. This means there is a notable decrease in F-Score across other varieties when the model is fine-tuned on a specific variety. This supports the idea that sarcasm classification requires the model to grasp linguistic and cultural nuances specific to a variety, making generalisation across varieties challenging. The corresponding results for bert are in Figure 6.
Cross-domain evaluation
Similarly, to assess the cross-domain robustness of our sentiment classification models, we conduct experiments with domain data from google, and reddit. Models trained on one domain are tested on the other domain, and vice-versa, for each of the three varieties of English. Figure 7 shows that models perform better when trained on in-domain data, for both mistral and bert. We observe that pre-trained mistral performs better for all varieties in comparison to results from the cross-domain experiments.
While mistral outperforms bert at both in- and cross-domain experiments, there is significant variation in cross-domain results, highlighting the challenge in generalization, given domain variance in fine-tuned data. Moreover, the direction of cross-domain experiments also affects performance. Both models perform better for reddit $\rightarrow$ google, suggesting that models may be better at transferring sentiment from a relatively informal domain. The overall difference in cross-domain performance is significantly higher for inner-circle varieties, compared to en-IN, indicating that fine-tuning with mixed domain data may be less effective for such varieties. Across varieties, however, en-UK, an inner-circle variety, performs better for google $\rightarrow$ reddit, whereas for the bert model, this is observed with the en-IN variety.
8. Error Analysis
We randomly sample 30 examples misclassified by mistral, across all domains and varieties. We hope that these error categories will be helpful for future researchers who use the BESSTIE dataset. For sentiment classification, we identify the following error categories:
  1. (1)
    Mixed Sentiment Mis-classification: The model struggles to classify text containing both positive and negative sentiment, as in ‘The outside pool has been completely redesigned and although looks good the depth of the pool has been downsized.’.
  2. (2)
    Insufficient Context: Consider the Reddit comment ‘Darwin for getting its own dot and big font on any map of the World.’. The sentiment can only be identified with the context/information from the previous exchanges in the thread.
For sarcasm classification, the error categories are:
  1. (1)
    Contextual Sarcasm: Some instances of sarcasm rely heavily on contextual or conversational cues that are not present in the text on its own. For example, in the Reddit comment, ‘Atleast Spiderman covered his face before answering.’, the sarcastic intent can only be understood within the context of the preceding discussion in the thread.
  2. (2)
    Implicit Sarcasm: Instances where sarcasm is expressed subtly, through tone or mild exaggeration, without explicit contradictions. The example, ‘Makes more sense to just bar his company from getting further contracts whilst he is running it.’, is misidentified as not sarcastic.
  3. (3)
    False Positives from Negative Connotations: Mildly negative statements are often misclassified as sarcastic due to their phrasing or sentiment. This error is identified for examples such as ‘Sorry gal, what you experienced pretty much IS the dating culture here.’.
9. Conclusion
Biases in LLMs have been reported for several language varieties. This paper presents BESSTIE, a benchmark for sentiment and sarcasm classification across three varieties of English: Australian (en-AU), Indian (en-IN), and British (en-UK). Our evaluation spans nine LLMs, six encoders and three decoders, assessed on datasets collected from Google Places reviews (google) and Reddit comments (reddit). The models perform consistently better on en-AU and en-UK (i.e., inner-circle varieties) than on en-IN (i.e., the outer-circle variety) for both tasks. Although models report high performance on sentiment classification (F-Score of 0.81 and 0.75, for google and reddit respectively), they struggle with detecting sarcasm (F-Score of 0.59 on reddit), indicating that sarcasm classification is still largely unsolved. This is supported by our cross-variety evaluation which reveals limited generalisation capability for sarcasm classification, where cultural and contextual understanding is crucial. Notably, monolingual models marginally outperform multilingual models, suggesting that language diversity in pre-training does not extend to varieties of a language. In conclusion, BESSTIE will provide a way to measure bias of LLMs towards non-standard English varieties for the two tasks. By highlighting gaps in the model performance, especially for (a) sarcasm classification in general; and (b) outer-circle varieties like en-IN for both the tasks, this benchmark will be useful to evaluate dialectal biases of language models, and contribute to research towards equitable performance across all varieties of English.
Acknowledgments
This work was funded by Google’s Research Scholar grant awarded in 2024.

### v3

@@
 BESSTIE: A Benchmark for Sentiment and Sarcasm Classification for Varieties of English
-Abstract.
-Despite large language models (LLMs) being known to exhibit bias against non-mainstream varieties, there are no known labeled datasets for sentiment analysis of English.
+Abstract
+Despite large language models (LLMs) being known to exhibit bias against non-standard language varieties, there are no known labelled datasets for sentiment analysis of English.
 To address this gap, we introduce besstie, a benchmark for sentiment and sarcasm classification for three varieties of English: Australian (en-AU), Indian (en-IN), and British (en-UK).
-Using web-based content from two domains, namely, Google Place reviews and Reddit comments, we collect datasets for these language varieties using two methods: location-based and topic-based filtering.
+We collect datasets for these language varieties using two methods: location-based for Google Places reviews, and topic-based filtering for Reddit comments.
+To assess whether the dataset accurately represents these varieties, we conduct two validation steps: (a) manual annotation of language varieties and (b) automatic language variety prediction.
 Native speakers of the language varieties manually annotate the datasets with sentiment and sarcasm labels.
+We perform an additional annotation exercise to validate the reliance of the annotated labels.
 Subsequently, we fine-tune nine large language models (LLMs) (representing a range of encoder/decoder and mono/multilingual models) on these datasets, and evaluate their performance on the two tasks.
-Our results reveal that the models consistently perform better on inner-circle varieties (i.e., en-AU and en-UK), with significant performance drops for en-IN, particularly in sarcasm detection.
+Our results show that the models consistently perform better on inner-circle varieties (i.e., en-AU and en-UK), in comparison with en-IN, particularly for sarcasm classification.
 We also report challenges in cross-variety generalisation, highlighting the need for language variety-specific datasets such as ours.
 besstie promises to be a useful evaluative benchmark for future research in equitable LLMs, specifically in terms of language varieties.
-The BESSTIE datasets, code, and models are currently available on request, while the paper is under review.
-Please email aditya.joshi@unsw.edu.au.
-bias, language variety, dialectal bias, language models, sentiment classification, sarcasm classification, benchmark, natural language processing
-ccs: Computing methodologies Natural language processingccs: Computing methodologies Language resources
-1. Introduction
-Benchmark-based evaluation (Socher et al., 2013) of large language models (LLMs) is the prevalent norm in natural language processing (NLP).
-Benchmarks provide labeled datasets (Wang et al., 2019; Muennighoff et al., 2023) for specific tasks so that LLMs can be evaluated against them.
-However, most NLP benchmarks do not contain text in language varieties such as national varieties, dialects, sociolects or creoles (Plank, 2022; Lent et al., 2024; Joshi et al., 2024).
-As a result, despite the superlative performance of LLMs on Standard American English, LLMs are not evaluated on other varieties of English.
-Knowing that LLMs can be biased against certain language varieties (as shown by  (Deas et al., 2023) for African-American English), little empirical evidence is found for their bias towards several language varieties, particularly those of English.
-An exception is Multi-VALUE (Ziems et al., 2023) which creates synthetic datasets of language varieties using linguistically-informed syntactic transformations and reports a degraded performance on the GLUE benchmark.
-However, such synthetic texts are insufficient in measuring language variety bias because language varieties are more than syntax and encompass orthography, vocabulary, and cultural pragmatics (Nguyen, 2021).
-As a result, labeled datasets of natural text are crucial to measure the bias in LLMs towards these varieties.
-Benchmark Sent. Sarc. Eng. Var.
-Cieliebak et al. (2017) (Cieliebak et al., 2017)
-Wang et al. (2018) (Wang et al., 2018)
-Alharbi et al. (2020) (Alharbi et al., 2020)
-Abu Farha et al. (2021) (Abu Farha et al., 2021)
-Elmadany et al. (2023) (Elmadany et al., 2023)
-Faisal et al. (2024) (Faisal et al., 2024)
+The besstie dataset is publicly available at: https://huggingface.co/datasets/unswnlporg/BESSTIE.
+BESSTIE: A Benchmark for Sentiment and Sarcasm Classification for Varieties of English
+Dipankar Srirag1 Aditya Joshi1 Jordan Painter2 Diptesh Kanojia2 1University of New South Wales, Sydney, Australia 2Institute for People-Centered AI, University of Surrey, Surrey, United Kingdom {d.srirag,aditya.joshi}@unsw.edu.au{jp1106,d.kanojia}@surrey.ac.uk
+1 Introduction
+Benchmark-based evaluation Socher et al.
+(2013) of large language models (LLMs) is the prevalent norm in natural language processing (NLP).
+Benchmarks provide labelled datasets Wang et al.
+(2019); Muennighoff et al.
+(2023) for specific tasks so that LLMs can be evaluated against them.
+However, most NLP benchmarks do not contain text in language varieties such as national varieties, dialects, sociolects or creoles Plank (2022); Lent et al.
+(2024); Joshi et al.
+(2025).
 […]
-Table 1.
-Comparison of BESSTIE with past benchmarks for sentiment or sarcasm classification.
+Table 1: Comparison of besstie with past benchmarks for sentiment or sarcasm classification.
 ‘Sent.’ indicates sentiment classification, ‘Sarc.’ denotes sarcasm classification, ‘Eng.’ denotes English, and ‘Var.’ denotes language varieties.
 A checkmark (✓) denotes the availability of a particular feature, while a cross (✕) indicates its absence.
-To address this gap, we introduce BESSTIE, A BEnchmark for Sentiment and Sarcasm classification for varieTIes of English, namely, Australian (en-AU), Indian (en-IN), and British (en-UK).
-BESSTIE comprises a manually labeled dataset of textual posts along with performance results of fine-tuned LLMs on the two classification tasks.
+Figure 1: Creating dataset for besstie; Collection in Section 2.1, Quality Assessment in Section 2.2, Annotation in 2.3.
+As a result, despite the superlative performance of LLMs on Standard American English, LLMs are not evaluated on other varieties of English.
+Knowing that LLMs can be biased against certain varieties of English, as shown by Deas et al.
+(2023) and Srirag et al.
+(2025) for African-American English and Indian English respectively, little empirical evidence is found for their bias towards several other varieties.
+An exception is Multi-VALUE Ziems et al.
+(2023) which creates synthetic datasets of language varieties using linguistically-informed syntactic transformations and reports a degraded performance on the GLUE benchmark.
+However, such synthetic texts do not accurately represent the spoken or written forms of the language variety because language varieties are more than syntax and encompass orthography, vocabulary, and cultural pragmatics Nguyen (2021).
+As a result, labelled datasets of natural text are crucial to measure the bias in LLMs towards these varieties.
+Acknowledging the role of the sentiment classification task Socher et al.
+(2013) in GLUE, we introduce besstie, A BEnchmark for Sentiment and Sarcasm classification for varieTIes of English, namely, Australian (en-AU), Indian (en-IN), and British (en-UK).
+The relationship between sentiment and sarcasm classification as related tasks is well-understood in the context of sentiment analysis Chauhan et al.
+(2020).
+besstie comprises a manually labelled dataset of textual posts along with performance results of fine-tuned LLMs on the two classification tasks.
 The dataset consists of text samples collected from two web-based domains, namely, Google Places reviews, and Reddit comments, using two filtering methods: location-based and topic-based filtering respectively.
+We validate the representation of these varieties through a two-step process: (a) a manual annotation exercise and (b) automatic variety identification using models fine-tuned on the ICE-Corpora Greenbaum and Nelson (1996).
 Following that, we obtain manual annotations for boolean sentiment and sarcasm labels from native speakers of the language varieties.
-The novelty of the BESSTIE dataset can be seen in Table 1.
-BESSTIE improves not only upon popular sentiment/sarcasm-labeled datasets but also a recent dialectal benchmark, DialectBench (Faisal et al., 2024), which contains dialectal datasets for sentiment classification but for languages other than English.
-Therefore, the BESSTIE dataset is novel in the introduction of new language varieties (specifically, varieties of English) and, an additional task, sarcasm classification.
-We then use the dataset to model both the tasks (sentiment and sarcasm classification) as binary classification problems.
+We then use the dataset to model both the tasks as binary classification problems.
 We evaluate nine fine-tuned LLMs, spanning encoder/decoder and monolingual/multilingual models.
@@
 The performance on sarcasm classification is consistently low across the three varieties, providing pointers for future work.
-Following that, we examine the aggregate performance for the three varieties, the two tasks and model attributes (encoder/decoder and monolingual/multilingual).
-We also examine cross-variety and cross-domain performance on all models highlighting the need for BESSTIE.
+Following that, we examine the aggregate performance for the three varieties, the two tasks, and model attributes (encoder/decoder and monolingual/multilingual).
+We also examine cross-variety performance on all models highlighting the need for besstie.
 besstie evaluation provides a solid baseline to examine biases of contemporary LLMs, and will help to develop new methods to mitigate these biases.
-The contributions of this work are:
-  • Creation of a manually annotated dataset of user-generated content with sentiment and sarcasm labels for three varieties of English, called BESSTIE.
-  • Evaluating nine LLMs on BESSTIE and analysing their performance to identify challenges arising due to language varieties.
-An overview of BESSTIE is described in Figure 1.
-The paper describes the dataset (Section 4; referring to ‘Collection’ and ‘Annotation’ in the figure) and the evaluation methodology (Section 5; ‘Evaluation’ in the Figure) of BESSTIE.
-The experiment setup and results are in Sections 6 and 7 respectively, followed by error analysis in Section 8.
-2. Related work
-NLP Benchmarks such as GLUE (Wang et al., 2018), SuperGLUE (Wang et al., 2019), and DynaBench (Kiela et al., 2021) have played a pivotal role in evaluating LLMs on multiple NLP tasks, including sentiment classification.
+The novelty of the besstie dataset can be seen in Table 1.
+besstie improves not only upon popular sentiment/sarcasm-labelled datasets but also a recent dialectal benchmark, DialectBench Faisal et al.
+(2024), which covers sentiment classification but for dialects of languages other than English.
+Therefore, the besstie dataset is novel in the introduction of new language varieties (specifically, varieties of English) for both sentiment and sarcasm classification.
+The contributions of this work are: (a) Creation of a manually annotated dataset of user-generated content with sentiment and sarcasm labels for three varieties of English, called besstie; (b) Evaluation of nine LLMs on besstie and analysing their performance to identify challenges arising due to language varieties.
+2 Dataset Creation
+Figure 1 shows the process followed to create a manually annotated dataset for sentiment and sarcasm classification.
+2.1 Data Collection
+We collect textual posts from two domains: Google Places reviews and Reddit comments, using two filtering methods: location-based and topic-based respectively.
+Location-based filtering implies that we select reviews that were posted for locations in the three countries.
+Specifically, we collect reviews posted in cities in AU, IN and UK, and their corresponding ratings ($1$ to $5$ stars, where $5$ is the highest.) using the Google Places API [^2] https://developers.google.com/maps/documentation/places/web-service/overview; Accessed on 10 February 2025 from the place types, also defined by the API.
+The criteria for city selection are based on population thresholds specific to each country (en-AU: $20$K; en-IN:$100$K; en-UK: $50$K).
+We filter out any non-English content using language probabilities calculated by fastText Grave et al.
+(2018) word vectors, using a threshold of $0.98$.
+Additionally, to mitigate the risk of including reviews written by tourists rather than residents, we exclude reviews from locations designated as ‘tourist attractions’ by the Google Places API.
+[…]
+Table 2: Performance of distil on sentiment classification with different labelling schemes.
+naive refers to labels based on 1 and 5 stars, while ours corresponds to labels based on 2 and 4 stars.
+$\mu$ denotes the average F-Score across all varieties
+We also conduct a preliminary investigation using DistilBERT-Base (distil;Sanh et al., 2020) to examine the influence of label semantics, the star ratings, on the performance of sentiment classification.
+We experiment with distil on the said task using the collected Google Places reviews under two distinct labelling schemes: (a) extreme ratings, i.e., 1 and 5 stars (naive); and (b) moderate ratings, i.e., 2 and 4 stars (ours).
+Table 2 presents the performance of distil for each variety under both schemes, reported using the macro-averaged F-Score.
+We observe that distil yields strong performance when trained on naive labels, achieving an average $\mu$ of 0.96 across the three varieties.
+However, task performance drops considerably with the use of the more ambiguous ours labels, highlighting that they introduce increased difficulty and nuanced content.
+Following these findings, we filter the dataset to retain only reviews with intermediate ratings (2 and 4 stars), to better capture this nuance and avoid model overfitting to well-distinguished and polarised examples.
+To create the reddit subset, we employ topic-based filtering and choose up to four subreddits per variety (en-AU: ‘melbourne’, ‘AustralianPolitics’,‘AskAnAustralian’; en-IN: ‘India’, ‘IndiaSpeaks’, ‘BollyBlindsNGossip’; en-UK: ‘England’, ‘Britain’, ‘UnitedKingdom’, ‘GreatBritishMemes’), where the topics are determined by native speakers of these language varieties.
+We select these subreddits based on the understanding that they feature discussions specific to a variety, making it highly likely that the post authors use the corresponding language variety.
+For each variety, we scrape $12,000$ comments evenly across the selected subreddits, capping at $20$ comments per post and focusing on recent posts.
+These are then randomly sampled and standardised to $3,000$ comments per variety before manual annotation.
+We discard information such as user identifiers and post identifiers to maintain the anonymity of the user.
+Figure 2: Confusion matrix showing the overlap in variety annotations between annotators, ant-IN, and ant-UK.
+Rows represent the labels assigned by ant-IN, and columns represent the labels assigned by ant-UK.
+The principle diagonal elements indicate agreement between annotators, while off-diagonal elements highlight disagreements.
+2.2 Quality Assessment
+While previous studies utilise location-based filtering Blodgett et al.
+(2016); Sun et al.
+(2023), we examine if the collected text from the two methods is indeed from these language varieties.
+To address this, we conduct two evaluations: (a) Manual annotation for national variety identification, and (b) Automated validation of national varieties.
+Manual annotation
+We request en-IN and en-UK annotators, asking them to manually identify the variety of a given text.
+Specifically, they label a random sample of $300$ texts ($150$ from reviews and $150$ from comments) as en-AU, en-IN, en-UK, or Cannot say.
+Annotator agreement is measured using Cohen’s kappa ($\kappa$).
+With respect to the true label (based on the location or topic), results show higher agreement with the en-IN annotator ($\kappa$ = 0.41) compared to the en-UK annotator ($\kappa$ = 0.34), indicating fair to moderate consistency.
+The inter-annotator agreement itself is $0.26$.
+Figure 2 shows that annotators find it difficult to agree on an inner-circle variety, and have the highest agreement when identifying the en-IN variety, with 46 agreements, demonstrating that the subset reliably represents this variety.
+As a result, the two subsets (reviews and comments) collectively form a good representative sample for language varieties.
+Variety Subset $P(\bar{\texttt{eng}})$ $P(\bar{\texttt{v}})$ F-Score en-AU google 0.99 0.99 0.99 reddit 0.98 0.95 0.93 en-IN google 0.99 0.94 0.91 reddit 0.87 0.78 0.69 en-UK google 0.99 0.99 0.99 reddit 0.98 0.93 0.90
+Table 3: Data quality for each variety and subset.
+$P(\bar{\texttt{eng}})$ is the average language probability (calculated using fastText Grave et al.
+(2018) word vectors) of a review or comment being in English.
+$P(\bar{\texttt{v}})$ represents the average variety probability of a review or comment being in the corresponding variety.
+F-Score measures the performance of the variety predictor.
+Automated validation
+We use two predictors, a language predictor, and a variety predictor, to perform an automatic evaluation of our dataset.
+For language predictor, we use fastText word vectors and extract the probability of the text being in English.
+We then fine-tune distil on ICE-Corpora, to use as a variety predictor.
+We model the task as binary classification (i.e., inner-circle vs.
+outer-circle variety classification).
+We use the ICE-Australia Smith and Peters (2023) and ICE-India subsets of the corpus [^3] Due to lack of access to ICE-GB, we do not include en-UK variety in training the predictor., focusing on the transcriptions from unprompted monologues and private dialogues (S1A and S2A headers).
+Similar to language predictor, we extract the probability of text being in the corresponding variety.
+$P(\bar{\texttt{eng}})$ in Table 3 is the probability that the text is written in English.
+$P(\bar{\texttt{v}})$ is the average variety probability.
+High values indicate that the variety predictor is capable of discerning text written in inner-circle varieties from those written in the outer-circle variety (i.e., en-IN).
+The lower language and variety probabilities on the reddit subset from en-IN, with 0.87 and 0.78 respectively, can be attributed to the code-mixed text in the subset.
+These results along with our manual annotation exercise show that the collected data represents the three language varieties.
+The probabilities from the model are in line with the ground truth, as measured using the F-Score.
+Variety Sent. Sarc. en-AU 0.61 0.47 en-IN 0.65 0.51 en-UK 0.79 0.63
+Table 4: Inter-annotator agreement for the validation annotation exercise.
+Agreement is measured between the original annotator and an independent annotator for each language variety.
+Here Sent.
+and Sarc.
+denote sentiment and sarcasm labels respectively.
+Variety Subset Train Valid Test % Pos.
+Sent.
+% Pos.
+Sarc.
+Avg.
+no.
+of words en-AU google 946 130 270 73% 7% 63.97 reddit 1763 241 501 32% 42% 51.72 en-IN google 1648 225 469 75% 1% 44.34 reddit 1686 230 479 25% 13% 26.92 en-UK google 1817 248 517 75% 0% 72.21 reddit 1007 138 287 12% 22% 38.04
+Table 5: Dataset statistics for each variety, subset, split, and label type. % Pos. Sent. and % Pos. Sarc. indicates the proportion of samples with positive sentiment and true sarcasm respectively.
+2.3 Annotation
+We hire one annotator [^4] Two of these are the authors of the paper who were also involved in the evaluation.
+each for the three language varieties.
+The annotators assign the processed reviews and posts two labels: sentiment and sarcasm.
+The choice of labels given are negative, positive, and discard.
+We instruct the annotators to use the discard label for uninformative examples with no apparent polarity and in a few cases, for computer-generated messages.
+The texts with discard label are discarded.
+We conduct an additional annotation exercise to evaluate the reliability of the sentiment and sarcasm annotations.
+For this, we randomly selected 50 instances from each variety, annotated by the original annotators, and hire a new independent annotator for each variety.
+The inter-annotator agreement between in the original and the independent annotators, measured by $\kappa$, are reported in the Table 4.
+The absolute agreement values for sarcasm, though lower than those for sentiment, are comparable to prior work.
+In particular, Joshi et al.
+(2016) report a $\kappa$ of 0.44 on their sarcasm labels.
+These scores suggest that both sentiment and sarcasm labels in our dataset are annotated with high degree of reliability.
+We compensate the annotators at the casual employment rate, equivalent to 22 USD per hour, as prescribed by the host institution.
+We provide the annotation guideline in Appendix A.
+2.4 Dataset Statistics
+The resultant besstie dataset contains annotations for two tasks: sentiment and sarcasm classification, for the two domains and three varieties.
+Table 5 provides a breakdown of the annotated dataset (consisting of google and reddit), including the number of samples in each split (Train, Validation, and Test), along with average word length.
+The % Pos.
+Sent.
+and % Pos.
+Sarc.
+indicates the proportion of samples with positive sentiment and true sarcasm respectively.
+These values are the same for the train, validation, and test set since we perform stratified sampling.
+We present example comments along with their corresponding sentiment and sarcasm labels in Appendix B.
+We also confirm that the text collected for Reddit subset of the besstie is generated in 2024.
+While this subset represents a static snapshot, its recency allows it to reflect contemporary linguistic usage effectively.
+Figure 3: Performance comparison of various models on the sentiment classification task across different English varieties (en-AU, en-IN, and en-UK).
+Figure 4: Performance comparison of various models on the sarcasm classification task across different English varieties (en-AU, en-IN, and en-UK).
+3 Experiment Details
+We conduct our experiments on a total of nine LLMs.
+These include six encoder models, three pre-trained on English corpora: BERT-Large (bertDevlin et al.
+(2019), RoBERTa-Large (robertaLiu et al.
+(2019), ALBERT-XXL-v2 (albertLan et al.
+(2020); and three multilingual models: Multilingual-BERT-Base (mbert), Multilingual-DistilBERT-Base (mdistilSanh et al.
+(2020), XLM-RoBERTa-Large (xlm-rGoyal et al.
+(2021).
+We also evaluate three decoder models, one pre-trained on English corpora: Gemma2-27B-Instruct (gemmaGemma Team et al.
+(2024) and two multilingual models: Mistral-Small-Instruct-2409 (mistralJiang et al.
+(2023), Qwen2.5-72B-Instruct (qwenYang et al.
+(2024).
+The model-specific details including architecture, language coverage of pre-training corpus, and number of parameters are described in Appendix C.
+While encoder models are fine-tuned with full precision, we fine-tune quantised decoder models using QLoRA Dettmers et al.
+(2023) adapters, targeting all linear layers.
+All models (including the variety predictor) are fined-tuned for 30 epochs, with a batch size of 8 and Adam optimiser.
+We choose an optimal learning rate by performing a grid search over the following values: 1e-5, 2e-5, and 3e-5.
+All experiments are performed using two NVIDIA A100 80GB GPUs.
+We fine-tune encoder models using cross-entropy loss, weighted on class distribution (positive: ‘1’ and negative: ‘0’ labels).
+In contrast, for decoder models, we perform zero-shot instruction fine-tuning Ouyang et al.
+(2022), using maximum likelihood estimation – a standard learning objective for causal language modeling Jain et al.
+(2023).
+For decoder models, we use two task-specific prompts, one for each task. For sentiment classification, we the prompt the model with:
+“Generate the sentiment of the given text. 1 for positive sentiment, and 0 for negative sentiment. Do not give an explanation.”
+Similarly, for sarcasm classification, we use:
+“Predict if the given text is sarcastic. 1 if the text is sarcastic, and 0 if the text is not sarcastic. Do not give an explanation.”
+Given the text and a task-specific prompt, the expected behaviour of the LLM is to generate either 1 (for positive) or 0 (for negative).
+The forms of these prompts are experimentally determined using a few test examples.
+Finally, we report our task performances on three macro-averaged F-Score, i.e., unweighted average, disregarding the imbalance in label distribution.
+4 Results
+We present our results and subsequent discussions to address the following questions: (a) How well do current LLMs perform on the benchmark tasks?
+(Section 4.1); (b) How do factors such as model properties and domain affect the task performance?
+(Section 4.2); (c) Can models trained on one language variety generalise well to others?
+(Section 4.3).
+4.1 Task Performance
+We first present our results on the two tasks where models are trained and evaluated on the same variety.
+Figures 3 and 4 describe the model performances, reported using F-Score, on the sentiment and sarcasm classification tasks respectively.
+While sentiment classification is performed using both google and reddit subsets, due to the absence of sarcasm labels in google subset (as shown in Table 5), sarcasm classification is conducted only on reddit subset.
+Table 6 shows that models, on average, perform better on both tasks when trained on en-AU variety, closely followed by performances on en-UK variety.
+Conversely, models consistently perform the worst on en-IN.
+Domain-Task en-AU en-IN en-UK
+google-Sentiment 0.94 0.64 0.86
+reddit-Sentiment 0.78 0.69 0.78
+reddit-Sarcasm 0.62 0.56 0.58
+[…]
+Table 6: Task performances, averaged across all models, on all varieties. $\mu$ represents the average performance across all domain-task pairs.
+Sentiment Classification
+Table 7 shows that, for sentiment classification, mistral achieves the highest average performance across all varieties, with an F-Score of 0.91 on the google subset and 0.84 on the reddit subset.
+In contrast, qwen reports the lowest average performance across all varieties on the google subset, with an average F-Score of 0.45, while gemma exhibits the lowest average performance across all varieties on the reddit subset, with an F-Score of 0.60.
+Sarcasm Classification
+Similarly, for sarcasm classification, both mistral and mbert achieve the highest average performance across all varieties, with an F-Score of 0.68, as shown in Table 7.
+In contrast, gemma again reports the lowest average performance across all varieties, with an average F-Score of 0.45.
+The lower performance on sarcasm classification as compared to sentiment classification is expected because sarcasm often relies on not just subtle linguistic cues but also localised contemporary contextual information Abercrombie and Hovy (2016).
+4.2 Impact of Models and Domain
+The average results for the models highlight the need to further probe into factors influencing model performances, namely, model properties and domain.
+Model properties
+The results ($\mu$ values) in the corresponding columns of Table 7 describe the model performances (reporting F-Score), grouped based on model properties (encoder/decoder, mono/multilingual).
+The values show a consistent trend where encoder models report higher performance than decoder models across different tasks.
+This performance gap is expected as encoder architecture is inherently better suited for sequence classification tasks, while decoder architecture is optimised for text generation tasks.
+We also find that, although marginal, monolingual models report a higher average performance compared to multilingual models.
+This performance gap is prominent on the task involving google subset, while multilingual models perform better on tasks involving reddit subset.
+google vs. reddit
+Our findings (the last row $\mu$ for the task columns in Table 7), indicate that models perform better on google subset, achieving an F-Score of 0.81, compared to an F-Score of 0.75 on reddit subset.
+This difference can be explained by the distinct writing styles of the two sources.
+Reviews in google are generally more formal and informative, while posts and comments from reddit often exhibit a short (as shown in Table 5), conversational tone typical of social media.
+Additionally, posts and comments in the reddit subset come from forums frequented by local speakers, which means the language and expressions used are more reflective of the variety and cultural nuances.
+We present results from cross-domain evaluation in the Appendix D.
+[…]
+Table 7: Model performances, averaged across all varieties, on sentiment and sarcasm classification tasks across two subsets (google and reddit).
+$\mu$ represents the average performance across all models.
+‘Enco.’ represents encoder models.
+‘Mono.’ represents monolingual models.
+Figure 5: Cross-variety performance analysis of mistral.
+The figure compares three different scenarios: pre-trained (PT), in-variety fine-tuning, and cross-variety fine-tuning for sentiment and sarcasm classification across all varieties.
+4.3 Cross-variety Evaluation
+Finally, we perform cross-variety evaluation.
+This refers to the scenario where the datasets used to train and test a model are from different groups (i.e., trained on en-AU, tested on en-IN, for example).
+We report the cross-variety results of our best-performing model, mistral.
+Figure 5 compares the performance of mistral across three variants: pre-trained, in-variety fine-tuning, and cross-variety fine-tuning, for the three dataset subsets using colour-coded matrices.
+The rows indicate the subset that the model is trained on (i.e., en-AU, en-IN, and en-UK), along with PT, indicating that the pre-trained mistral is used.
+Similarly, the columns indicate the test subset (i.e., en-AU, en-IN, and en-UK).
+The pre-trained variant of the model (the first row in each matrix) achieves high F-Score for en-AU and en-UK, for the sentiment classification task.
+Fine-tuning the model improves in-variety performance (shown by the principal diagonal of the matrices), with noticeable gains in F-Score on both subsets.
+The model performance on cross-variety evaluation remains relatively stable across different varieties of English, indicating that domain-specific data plays a minor role in influencing cross-variety generalisation for sentiment classification.
+The model performs significantly worse on sarcasm classification compared to sentiment classification.
+While in-variety evaluation shows that fine-tuning improves over the pre-trained model with substantial increases in F-Score, it negatively impacts cross-variety generalisation.
+This means there is a notable decrease in F-Score across other varieties when the model is fine-tuned on a specific variety.
+This supports the idea that sarcasm classification requires the model to grasp linguistic and cultural nuances specific to a variety, making generalisation across varieties challenging.
+Variety Sample dial coll cont code en-AU 70 9 28 6 - en-IN 90 97 33 3 8 en-UK 53 7 15 5 -
+Table 8: Counts of identified features in misclassified examples from mistral for each variety.
+dial denotes dialect features, coll denotes locale-specific colloquial expressions, cont denotes instances requiring additional context, and code denotes occurrences of code-mixed text.
+5 Error Analysis
+We randomly sample up to 30 examples misclassified by mistral, from each variety and domain-task configuration.
+We then manually analyse these misclassified examples to identify pervasive or obligatory dialect features, as defined by eWAVE Kortmann et al.
+(2020).
+A dialect feature is defined as any lexical or syntactic variation from standard English.
+In addition, we identify other features, including locale-specific colloquial expressions, instances that require additional context, and occurrences of code-mixed text.
+Table 8 summarises the sample sizes and the counts of these identified features for each variety.
+Detailed examples for each feature type are provided in Appendix E.
+The en-IN variety exhibits a high count of dialect features (97) relative to its sample size (90), suggesting that regional variations significantly challenge model performance.
+Locale-specific colloquial expressions are prevalent in all varieties (en-AU: 28; en-IN: 33; en-UK: 15), highlighting the impact of regional lexicons on model performance.
+These findings motivate future research in sentiment and sarcasm classification to better accommodate such colloquial expressions for inner-circle varieties, and to additionally adapt to dialect features for outer-circle varieties of English.
+6 Related work
+NLP Benchmarks such as GLUE Wang et al.
+(2018), SuperGLUE Wang et al.
+(2019), and DynaBench Kiela et al.
+(2021) have played a pivotal role in evaluating LLMs on multiple NLP tasks, including sentiment classification.
 However, these benchmarks are limited in their ability to capture the linguistic nuances of non-standard language varieties.
-Efforts to address this include datasets for African languages and dialects (Muhammad et al., 2023), Arabic dialects (Elmadany et al., 2023), and CreoleVal (Lent et al., 2023) which provides datasets for 28 creole languages.
-More recently, DIALECTBENCH (Faisal et al., 2024) is a collection of dialectal datasets, and associated NLP models across 281 dialectal varieties for several tasks.
+Efforts to address this include datasets for African languages and dialects Muhammad et al.
+(2023), Arabic dialects Elmadany et al.
+(2023), and CreoleVal Lent et al.
+(2024) which provides datasets for 28 creole languages.
+More recent progress includes DialectBench Faisal et al.
+(2024), which presents a collection of dialectal datasets and associated NLP models across 281 dialectal varieties for several tasks.
 However, the benchmark does not contain English dialectal datasets for sentiment and sarcasm classification.
-BESSTIE is the first benchmark for sentiment and sarcasm classification specifically for non-standard varieties of English.
-Similarly, several sarcasm classification datasets for standard variety of English exist, consisting of reddit comments (Khodak et al., 2018), amazon reviews (Filatova, 2012), and tweets (Ptáček et al., 2014; Painter et al., 2022; Abercrombie and Hovy, 2016; Oprea and Magdy, 2020).
-However, there is a notable absence of sarcasm classification datasets that account for non-standard varieties of English.
-ArSarcasm-v2 (Abu Farha et al., 2021) is a dataset annotated for sarcasm, sentiment, and the dialect of each tweet for Arabic.
-Abu Farha and Magdy (2021) benchmarked Arabic-specific transformer models on this dataset, noting that models trained on Arabic data, including varieties of the language, tend to perform better.
-We emphasise that the papers listed above approach benchmark creation in a similar manner as BESSTIE: they create an annotated dataset for a task, and test a broad range of models to highlight the potential for future work.
-In that sense, these benchmark papers acknowledge that creating a dataset and measuring bias is an important task achieved by papers that introduce benchmarks, without necessarily presenting a new method itself for the task.
-This holds true for our paper as well.
-3. Ethical Considerations
-The project received ethics approval from the Human Research Ethics Committee at UNSW, Sydney (reference number: iRECS6514).
-We treat the reviews as separate texts and do not aggregate them by user in any way.
+Similarly, several sarcasm classification datasets for standard variety of English exist, consisting of reddit comments Khodak et al.
+(2018), amazon reviews Filatova (2012), and tweets Ptáček et al.
+(2014); Painter et al.
+(2022); Abercrombie and Hovy (2016); Oprea and Magdy (2020).
+While datasets like ArSarcasm-v2 Abu Farha et al.
+(2021) present annotations for sarcasm, sentiment, and the dialect of each tweet for Arabic, there is a notable absence of sarcasm classification datasets that account for non-standard varieties of English.
+Similar to their work, we also create a dataset that contains both sentiment and sarcasm labels.
+To the best of our knowledge, besstie is the first benchmark for sentiment and sarcasm classification specifically for varieties of English.
+7 Conclusion
+This paper presents besstie, a benchmark for sentiment and sarcasm classification across three varieties of English: Australian (en-AU), Indian (en-IN), and British (en-UK).
+Our evaluation spans nine LLMs, six encoders and three decoders, assessed on datasets collected from Google Places reviews (google) and Reddit comments (reddit).
+The models perform consistently better on en-AU and en-UK (i.e., inner-circle varieties) than on en-IN (i.e., the outer-circle variety) for both tasks.
+We verify the variety representation of collected data using: (a) manual annotation and (b) automated validation.
+We also perform an additional annotation exercise to verify the reliability of the annotations.
+Although models report high performance on sentiment classification (F-Score of 0.81 and 0.75, for google and reddit respectively), they struggle with detecting sarcasm (F-Score of 0.59 on reddit), indicating that sarcasm classification is still largely unsolved.
+This is supported by our cross-variety evaluation which reveals limited generalisation capability for sarcasm classification, where cultural and contextual understanding is crucial.
+Notably, monolingual models marginally outperform multilingual models, suggesting that language diversity in pre-training does not extend to varieties of a language.
+Our error analysis provides potential considerations to drive future research in sentiment and sarcasm classification to better accommodate colloquial expressions for inner-circle varieties, and dialect features for outer-circle varieties of English.
+In conclusion, besstie provides a resource to measure bias in LLMs towards non-standard English varieties for the two tasks.
+Limitations
+Our assumption of national varieties as representative forms of dialect is a simplification.
+Within each locale, there are significant regional, sociolectal, and generational differences.
+For example, Australian English spoken in Sydney may differ significantly from that in Perth.
+The language used in online communications evolves rapidly, driven by cultural trends, viral content, and changing societal norms.
+Phrases, slang, and expressions can gain popularity and fade quickly.
+A static dataset, such as ours, is hence limited in reflecting these continuous and dynamic changes in language use.
+While our efforts to remove non-English text were thorough, there may still be instances where language mixing or code-switching occurs, especially in the Indian English subset, affecting model performance.
+The annotation process was limited by the number of human annotators and their subjective interpretations.
+While we perform an additional annotation exercise to validate the annotation process, use of a single annotator can introduce individual biases.
+We also note that comments from Reddit may differ from other social media sites, such as Twitter/X, in terms of user demographics and discourse style.
+This potentially influences the nature of topic-based discussions captured in our dataset.
+Our reliance on Reddit data is due to cost-related constraints associated with accessing Twitter/X API.
+Ethical Considerations
+The project received ethics approval from the Human Research Ethics Committee at UNSW, Sydney (reference number: iRECS6514), the host organisation for this research.We treat the reviews as separate texts and do not aggregate them by user in any way.
 The Reddit comments and Google reviews used in this paper are from publicly available posts, accessible through official APIs.
@@
 The dataset, including sentiment and sarcasm annotations, is shared in a manner that preserves user privacy and abides by the rules and guidelines of the source platforms.
-Figure 1. Methodology of BESSTIE. (a) Collection is detailed in Section 4.1, (b) Annotation is detailed in 4.2, and (c) Evaluation is detailed in Section 5.
-4. Dataset Creation
-We now describe parts (a) and (b) of Figure 1 which result in a manually annotated dataset for sentiment and sarcasm classification.
-4.1. Data Collection
-We collect textual posts from two domains: Google Places reviews and Reddit comments, using two filtering methods: location-based and topic-based respectively.
-Location-based filtering implies that we select reviews that were posted for locations in the three countries.
-Specifically, we collect reviews posted in cities in AU, IN and UK, and their corresponding ratings ($1$ to $5$ stars, where $5$ is the highest.) using the Google Places API [^2] https://developers.google.com/maps/documentation/places/web-service/overview; Accessed on 23 July 2024 from the place types, also defined by the API.
-The criteria for city selection are based on population thresholds specific to each country (en-AU: $20$K; en-IN:$100$K; en-UK: $50$K).
-We further filter the dataset to only contain reviews with closer ratings ($2$ and $4$ stars).
-We also filter out any non-English content using language probabilities calculated by fastText (Grave et al., 2018) word vectors, using a threshold of $0.98$.
-Additionally, to mitigate the risk of including reviews written by tourists rather than residents, we exclude reviews from locations designated as ‘tourist attractions’ [^3] https://developers.google.com/maps/documentation/places/web-service/supported_types; Accessed on 13 November 2024.
-by the Google Places API.
-To create the reddit subset, we employ topic-based filtering and choose up to four subreddits per variety (en-AU: ‘melbourne’, ‘AustralianPolitics’,‘AskAnAustralian’; en-IN: ‘India’, ‘IndiaSpeaks’, ‘BollyBlindsNGossip’; en-UK: ‘England’, ‘Britain’, ‘UnitedKingdom’, ‘GreatBritishMemes’).
-We select these subreddits based on the understanding that they feature popular, discussions specific to a variety, making it highly likely that contributors use the appropriate language variety.
-For each variety, we scrape $12,000$ comments evenly across the selected subreddits, capping at $20$ comments per post and focusing on recent posts.
-These are then randomly sampled and standardized to $3,000$ comments per variety before manual annotation.
-4.2. Annotation
-We hire three annotators [^4] Two of these are the authors of the paper who were also involved in the evaluation.
-each for the three language varieties.
-The annotators assign the processed reviews and posts two labels: sentiment and sarcasm.
-The choice of labels given are negative, positive, and discard.
-We instruct the annotators to use the discard label for uninformative examples with no apparent polarity and in a few cases, for computer-generated messages.
-We compensate the annotators at the casual employment rate prescribed by the host institution.
-Figure 2.
-Confusion matrix showing the overlap in variety annotations between annotators, ant-IN, and ant-UK.
-Rows represent the labels assigned by ant-IN, and columns represent the labels assigned by ant-UK.
-The principle diagonal elements indicate agreement between annotators, while off-diagonal elements highlight disagreements.
-4.3. Data Quality & Statistics
-Data Quality
-While previous studies utilise location-based filtering to evaluate the robustness of language technologies across varieties (Blodgett et al., 2016; Sun et al., 2023), it needs to be examined if the collected text is indeed from these national varieties.
-To address this, we conduct an annotation exercise with en-IN and en-UK annotators, asking them to manually identify the variety of a given text.
-Specifically, they label a random sample of $300$ texts ($150$ from reviews and $150$ from comments) as en-AU, en-IN, en-UK, or Cannot say.
-Annotator agreement is measured using Cohen’s kappa ($\kappa$).
-With respect to the true label (based on the location or topic), results show higher agreement with the en-IN annotator ($\kappa$ = 0.41) compared to the en-UK annotator ($\kappa$ = 0.34), indicating fair to moderate consistency.
-The inter-annotator agreement itself is $0.26$.
-Figure 2 shows that annotators find it difficult to agree on inner-circle varieties, and have the highest agreement when identifying en-IN variety, with 46 agreements.
-While agreement between annotators is limited for inner-circle varieties, the higher agreement for the en-IN indicates that the subset reliably represents this variety, thereby validating our collection methodology.
-As a result, the two subsets (reviews and comments) collectively form a good representative sample for language varieties.
-Note: We also attempt to fine-tune an encoder model fine-tuned on an unlabeled dialectal corpus, to predict the language varieties.
-However, this was unsuccessful due to: (a) unavailability of the dataset for en-UK, (b)the dataset being dated and inaccurate in the domain of the collected data.
+Acknowledgements
+This research was funded by Google’s Research Scholar grant, awarded to Aditya Joshi and Diptesh Kanojia in 2024.
+Appendix A Annotation Guidelines
+Task Overview
+Your task involves determining whether each Google review or Reddit comment expresses positive or negative sentiment and whether it uses sarcasm.
+Annotation Instructions
+For Sentiment Classification:
+  • Positive Sentiment (Label 1): Annotate a comment as positive if it reflects favorable emotions such as happiness, satisfaction, agreement, or excitement.
+  • Negative Sentiment (Label 0): Annotate a comment as negative if it conveys unfavorable emotions such as sadness, disappointment, frustration, or criticism.
+  • Neither Positive nor Negative (Discarded, Label 2): discard the comment if it does not convey clear positive or negative sentiment (mark the same sarcasm label as 2).
+For Sarcasm Classification:
+  • Sarcastic Comment (Label 1): Annotate if the comment uses irony or mockery to express contempt or ridicule.
+  • Non-Sarcastic Comment (Label 0): Annotate if the comment is straightforward and does not employ sarcasm or irony.
+Additional Guidelines
+Consider the tone and context of the comment when assigning sentiment or sarcasm labels.
+For sarcasm, look for indicators such as exaggerated language, contradictions, or unexpected statements.
+Base your decision on the sentiment explicitly expressed in the text, avoiding personal biases.
+It is acceptable to mark a comment as undecided or seek clarification if you cannot confidently assign a sentiment label.
+Appendix B Dataset Examples
 Variety Comment Sentiment Sarcasm
@@
 en-UK What a brave potatriot 0 1
-Table 2. Reviews and comments from different varieties with their corresponding annotated sentiment and sarcasm labels.
-Dataset Statistics
-The benchmark dataset contains annotations for two tasks: sentiment and sarcasm classification, distributed across the three varieties.
-Table 3 provides a breakdown of the annotated dataset (consisting of google and reddit), including the number of samples in each split (Train, Validation, and Test), along with average word length and language probability.
-The % Positive sentiment and % Positive sarcasm indicate the proportion of samples with positive sentiment and true sarcasm respectively.
-These values are the same for train, validation and test set since we perform stratified sampling.
-We present a few example comments along with their corresponding sentiment and sarcasm labels in Table 2.
-The examples illustrate the diversity of language use and the challenges involved in detecting both sentiment and sarcasm across the varieties.
-Variety Subset Train Valid Test % Positive Sentiment % Positive Sarcasm $P(\bar{\texttt{eng}})$ $\bar{w}$
-en-AU google 946 130 270 73% 7% 0.99 63.97
-reddit 1763 241 501 32% 42% 0.98 51.72
-en-IN google 1648 225 469 75% 1% 0.99 44.34
-reddit 1686 230 479 25% 13% 0.87 26.92
-en-UK google 1817 248 517 75% 0% 0.99 72.21
-reddit 1007 138 287 12% 22% 0.98 38.04
-Table 3.
-Dataset statistics for each variety, subset, split, and label type.
-$\mu$ represents the average probability of English.
-$P(\bar{\texttt{eng}})$ is the average language probability (calculated using fastText (Grave et al., 2018) word vectors) of a review or comment being in English.
-$\bar{w}$ denotes average number of words in a review/comment.
-5. Evaluation Methodology
-With the annotated dataset in place, we now discuss the evaluation methodology used to evaluate the LLMs on the dataset.
-The training strategies for encoder and decoder models are as follows. We train encoder models using cross-entropy loss, weighted on class distribution:
-$\mathcal{L}_{\texttt{{Encoder}}}=-\frac{1}{N}\sum_{j=1}^{N}\sum_{k=1}^{C}w_{k}\cdot y_{j,k}\cdot\log p_{j,k}$
-Here, $N$ denotes the total number of samples in the batch, $C$ represents the number of classes (2 since the tasks are binary classification), and $w_{k}$ is the weight assigned to class k, calculated based on the class distribution ($w_{\texttt{neg}}$ for the negative and $w_{\texttt{pos}}$ for the positive class).
-$y_{j}$ is the expected output label for the $j$th instance where $y_{j,1}=1$ if the label is 1 and $y_{j,0}=1$ if the label is 0.
-$p_{j,k}$ is the probability of the $j$th instance for $k$th class.
-In contrast, for decoder models, we perform zero-shot instruction fine-tuning (Ouyang et al., 2022), using maximum likelihood estimation – a standard learning objective for causal language modeling (Jain et al., 2023):
-$\mathcal{L}_{\texttt{{Decoder}}}=-\frac{1}{N}\sum_{j=1}^{N}\left\{\sum_{i=|\mathbf{x}^{j}|+1}^{|\mathbf{x}^{j}|+|\mathbf{t}^{j}|}\log p(\mathbf{x}_{i}^{j}|\mathbf{x}_{<i}^{j})\right\}$
-Here, $\mathbf{x}_{<i}^{j}=[\mathbf{x}_{1}^{j},\ldots,\mathbf{x}_{i-1}^{j}]$ denotes the subsequence before $\mathbf{x}_{i}^{j}$ and ||$\cdot$|| ⋅ | is the number of tokens.
-Specifically, $|t_{j}|$ is the number of generated tokens and is set to $1$.
-In other words, given the text and a task-specific prompt (given in Table 4555The forms are experimentally determined using a few test examples.), the expected behaviour of the LLM is to generate either 1 (for positive) or 0 (for negative).
-Task Prompt
-Sentiment Classification “Generate the sentiment of the given text. Do not give an explanation.”
-Sarcasm Classification “Predict if the given text is sarcastic. Do not give an explanation.”
-Table 4. Task-specific prompts used to fine-tune decoder models. Both prompts restrict the model from generating an explanation.
-6. Experiment Details
-We conduct our experiments on a total of nine LLMs.
-These include six encoder models, three trained on English corpora: BERT-Large (bert(Devlin et al., 2019), RoBERTa-Large (roberta(Liu et al., 2019), ALBERT-XXL-v2 (albert(Lan et al., 2020); and three multilingual models: Multilingual-BERT-Base (mbert), Multilingual-DistilBERT-Base (mdistil(Sanh et al., 2020), XLM-RoBERTa-Large (xlm-r(Goyal et al., 2021).
-We also evaluate three decoder models, one trained on English corpora: Gemma2-27B-Instruct (gemma(Gemma Team et al., 2024) and two multilingual models: Mistral-Small-Instruct-2409 (mistral(Jiang et al., 2023), Qwen2.5-72B-Instruct (qwen(Yang et al., 2024).
-The model-specific details including architecture, language coverage of pre-training corpus, and number of parameters are described in Table 5.
-While encoder models are fine-tuned with full precision, we fine-tune quantized decoder models using QLoRA (Dettmers et al., 2023) adapters, targeting all linear layers.
-All models are fined-tuned for 30 epochs, with a batch size of 8 and Adam optimiser.
-We choose an optimal learning rate by performing a grid search over the following values: 1e-5, 2e-5, and 3e-5.
-All experiments are performed using two NVIDIA A100 80GB GPUs.
-We report our task performances on three macro-averaged F-Score, i.e., unweighted average, disregarding the imbalance in label distribution.
+Table 9: Reviews and comments from different varieties with their corresponding annotated sentiment and sarcasm labels.
+The examples presented in Table 9 illustrate the diversity of language use and the challenges involved in detecting both sentiment and sarcasm across the varieties.
+Appendix C Additional Model Details
+We describe additional model details, including architecture, the language coverage in their pre-training data, and the number of trainable parameters in Table 10.
 Model Arch. Lang. # Params
@@
 xlm-r Encoder Multilingual 355 M
-[…]
+gemma Decoder English 27 B
 mistral Decoder Multilingual 22 B
 qwen Decoder Multilingual 72 B
-Table 5. Details of models including architecture (Arch.), language coverage in pretraining data (Lang.), and number of trainable parameters (# Params). Here, M is millions and B is billion.
-7. Results
-Figure 3. Performance comparison of various models on the sentiment classification task across different English varieties (en-AU, en-IN, and en-UK).
-7.1. Overall results
-We first present our results on the two tasks where models are trained and evaluated on the same variety.
-Figure 3 and 4 describe the model performances, reported using F-Score, on the sentiment and sarcasm classification tasks respectively.
-While sentiment classification is performed using both google and reddit subsets, due to the absence of sarcasm labels in google subset (as shown in Table 3), sarcasm classification is conducted only on reddit subset.
-Sentiment Classification
-Table 6 shows that, for sentiment classification, mistral achieves the highest average performance across all varieties, with an F-Score of 0.91 on the google subset and 0.84 on the reddit subset.
-In contrast, qwen reports the lowest average performance across all varieties on the google subset, with an average F-Score of 0.45, while gemma exhibits the lowest average performance across all varieties on the reddit subset, with an F-Score of 0.60.
-Figure 4. Performance comparison of various models on the sarcasm classification task across different English varieties (en-AU, en-IN, and en-UK).
-Sarcasm Classification
-Similarly, Table 7 shows that, for sarcasm classification, both mistral and mbert achieve the highest average performance across all varieties, with an F-Score of 0.68.
-In contrast, gemma again reports the lowest average performance across all varieties, with an average F-Score of 0.45.
-The lower performance on sarcasm classification as compared to sentiment classification is expected because sarcasm often relies on not just subtle linguistic cues but also localised contemporary contextual information (Abercrombie and Hovy, 2016).
-Model Enco. Mono. google-Sentiment reddit-Sentiment reddit-Sarcasm
-albert 0.88 0.75 0.65
-bert 0.88 0.80 0.60
-roberta 0.89 0.79 0.55
-mbert 0.88 0.65 0.68
-mdistil 0.86 0.78 0.51
-xlm-r 0.77 0.73 0.62
-gemma 0.79 0.60 0.45
-mistral 0.91 0.84 0.68
-qwen 0.45 0.82 0.52
-[…]
-Table 6.
-Model performances, averaged across all varieties, on sentiment and sarcasm classification tasks across two subsets (google and reddit).
-$\mu$ represents the average performance across all models.
-‘Enco.’ represents encoder models.
-‘Mono.’ represents monolingual models.
-7.2. Impact of Domain and Models
-The average results for the models highlight the need to further probe into factors influencing model performances, namely, domain, model properties and language varieties.
-The results in Tables 7 and 8 describe the model performances (reporting F-Score), grouped based on model properties.
-Finally, task performances, averaged across all models, on varieties are presented in Table 9.
-google vs. reddit
-Our findings (the last row $\mu$ in Table 6), indicate that models perform better on google subset, achieving an F-Score of 0.81, compared to an F-Score of 0.75 on reddit subset.
-This difference can be explained by the distinct writing styles of the two sources.
-Reviews in google are generally more formal and informative, while posts and comments from reddit often exhibit a short (as shown in Table 3), conversational tone typical of social media.
-Additionally, posts and comments in reddit come from forums frequented by local speakers, which means the language and expressions used are more reflective of the variety and cultural nuances.
-Domain-Task Enco. Deco.
-google-Sentiment 0.86 0.72
-reddit-Sentiment 0.75 0.75
-reddit-Sarcasm 0.60 0.55
-[…]
-Table 7.
-Task performances, averaged across all varieties, grouped based on model architectures.
-$\mu$ represents the average performance across all domain-task pairs.‘Enco.’ represents encoder models while ‘Deco.’ represents decoder models.
-Domain-Task Mono. Mult.
-google-Sentiment 0.86 0.78
-reddit-Sentiment 0.73 0.76
-reddit-Sarcasm 0.56 0.60
-[…]
-Table 8.
-Task performances, averaged across all varieties, grouped based on language diversity in pre-training data.
-‘Mono.’ represents monolingual models, and ‘Mult.’ represents multilingual models.
-$\mu$ represents the average performance across all domain-task pairs.
-Model properties
-Table 7 shows a consistent trend where encoder models report higher performance than decoder models across different tasks.
-This performance gap is expected as encoder architecture is inherently better suited for sequence classification tasks, while decoder architecture is optimised for text generation tasks.
-Table 8 shows that, although marginal, monolingual models report a higher average performance compared to multilingual models.
-This performance gap is prominent on the task involving google subset, while multilingual models perform better on tasks involving reddit subset.
-Domain-Task en-AU en-IN en-UK
-google-Sentiment 0.94 0.64 0.86
-reddit-Sentiment 0.78 0.69 0.78
-reddit-Sarcasm 0.62 0.56 0.58
-[…]
-Table 9. Task performances, averaged across all models, on all varieties. $\mu$ represents the average performance across all domain-task pairs.
-Comparison of language varieties
-Finally, Table 9 shows that models, on average, perform better on both tasks when trained on en-AU variety, closely followed by performances on en-UK variety.
-Both en-AU and en-UK are inner-circle varieties.
-Conversely, models consistently perform the worst on en-IN, an outer-circle variety.
-Figure 5.
-Cross-variety performance analysis of mistral.
-The figure compares three different scenarios: pre-trained (PT), in-variety fine-tuning, and cross-variety fine-tuning for sentiment and sarcasm classification across all varieties.
-Figure 6.
-Cross-variety performance analysis of bert.
-The figure compares two different scenarios: in-variety fine-tuning, and cross-variety fine-tuning for sentiment and sarcasm classification across all varieties.
-Figure 7. Cross-domain performance analysis of mistral and bert comparing a decoder and an encoder model.
-7.3. Cross-variety/domain evaluation
-Finally, we perform cross-variety/cross-domain evaluation.
-This refers to the scenario where the datasets used to train and test a model are from different groups (i.e., trained on en-AU, tested on en-IN, for example).
-The following discussion is focused on generalisation: “can models trained on one language variety generalise well to others”.
-Cross-variety evaluation
-We report the cross-variety results of our best-performing model, mistral.
-Figure 5 compares the performance of mistral across three variants: pre-trained, in-variety fine-tuning, and cross-variety fine-tuning, for the three dataset subsets using colour-coded matrices.
-The rows indicate the subset that the model is trained on (i.e., en-AU, en-IN, and en-UK), along with PT, indicating that the pre-trained mistral is used.
-Similarly, the columns indicate the test subset (i.e., en-AU, en-IN, and en-UK).
-The pre-trained variant of the model (the first row in each matrix) achieves high F-Score for en-AU and en-UK, for the sentiment classification task.
-Fine-tuning the model improves in-variety performance (shown by the principal diagonal of the matrices), with noticeable gains in F-Score on both subsets.
-The model performance on cross-variety evaluation remains relatively stable across different varieties of English, indicating that domain-specific data plays a minor role in influencing cross-variety generalisation for sentiment classification.
-As seen earlier, the model performs significantly worse on sarcasm classification compared to sentiment classification.
-While in-variety evaluation shows that fine-tuning improves over the pre-trained model with substantial increases in F-Score, it negatively impacts cross-variety generalization.
-This means there is a notable decrease in F-Score across other varieties when the model is fine-tuned on a specific variety.
-This supports the idea that sarcasm classification requires the model to grasp linguistic and cultural nuances specific to a variety, making generalisation across varieties challenging.
-The corresponding results for bert are in Figure 6.
-Cross-domain evaluation
-Similarly, to assess the cross-domain robustness of our sentiment classification models, we conduct experiments with domain data from google, and reddit.
+Table 10: Details of models including architecture (Arch.), language coverage in pre-training data (Lang.), and number of trainable parameters (# Params). Here, M is millions and B is billion.
+Appendix D Cross-domain Evaluation
+We conduct experiments with domain data from google, and reddit, to assess the cross-domain robustness of our sentiment classification models.
 Models trained on one domain are tested on the other domain, and vice-versa, for each of the three varieties of English.
@@
 We observe that pre-trained mistral performs better for all varieties in comparison to results from the cross-domain experiments.
+Figure 6: Cross-variety performance analysis of bert.
+The figure compares two different scenarios: in-variety fine-tuning, and cross-variety fine-tuning for sentiment and sarcasm classification across all varieties.
+The corresponding results for bert are in Figure 6.
+Figure 7: Cross-domain performance analysis of mistral and bert comparing a decoder and an encoder model.
 While mistral outperforms bert at both in- and cross-domain experiments, there is significant variation in cross-domain results, highlighting the challenge in generalization, given domain variance in fine-tuned data.
@@
 Both models perform better for reddit $\rightarrow$ google, suggesting that models may be better at transferring sentiment from a relatively informal domain.
-The overall difference in cross-domain performance is significantly higher for inner-circle varieties, compared to en-IN, indicating that fine-tuning with mixed domain data may be less effective for such varieties.
-Across varieties, however, en-UK, an inner-circle variety, performs better for google $\rightarrow$ reddit, whereas for the bert model, this is observed with the en-IN variety.
-8. Error Analysis
-We randomly sample 30 examples misclassified by mistral, across all domains and varieties.
-We hope that these error categories will be helpful for future researchers who use the BESSTIE dataset.
-For sentiment classification, we identify the following error categories:
-  1. (1)
-    Mixed Sentiment Mis-classification: The model struggles to classify text containing both positive and negative sentiment, as in ‘The outside pool has been completely redesigned and although looks good the depth of the pool has been downsized.’.
-  2. (2)
-    Insufficient Context: Consider the Reddit comment ‘Darwin for getting its own dot and big font on any map of the World.’.
-The sentiment can only be identified with the context/information from the previous exchanges in the thread.
-For sarcasm classification, the error categories are:
-  1. (1)
-    Contextual Sarcasm: Some instances of sarcasm rely heavily on contextual or conversational cues that are not present in the text on its own.
-For example, in the Reddit comment, ‘Atleast Spiderman covered his face before answering.’, the sarcastic intent can only be understood within the context of the preceding discussion in the thread.
-  2. (2)
-    Implicit Sarcasm: Instances where sarcasm is expressed subtly, through tone or mild exaggeration, without explicit contradictions.
-The example, ‘Makes more sense to just bar his company from getting further contracts whilst he is running it.’, is misidentified as not sarcastic.
-  3. (3)
-    False Positives from Negative Connotations: Mildly negative statements are often misclassified as sarcastic due to their phrasing or sentiment.
-This error is identified for examples such as ‘Sorry gal, what you experienced pretty much IS the dating culture here.’.
-9. Conclusion
-Biases in LLMs have been reported for several language varieties.
-This paper presents BESSTIE, a benchmark for sentiment and sarcasm classification across three varieties of English: Australian (en-AU), Indian (en-IN), and British (en-UK).
-Our evaluation spans nine LLMs, six encoders and three decoders, assessed on datasets collected from Google Places reviews (google) and Reddit comments (reddit).
-The models perform consistently better on en-AU and en-UK (i.e., inner-circle varieties) than on en-IN (i.e., the outer-circle variety) for both tasks.
-Although models report high performance on sentiment classification (F-Score of 0.81 and 0.75, for google and reddit respectively), they struggle with detecting sarcasm (F-Score of 0.59 on reddit), indicating that sarcasm classification is still largely unsolved.
-This is supported by our cross-variety evaluation which reveals limited generalisation capability for sarcasm classification, where cultural and contextual understanding is crucial.
-Notably, monolingual models marginally outperform multilingual models, suggesting that language diversity in pre-training does not extend to varieties of a language.
-In conclusion, BESSTIE will provide a way to measure bias of LLMs towards non-standard English varieties for the two tasks.
-By highlighting gaps in the model performance, especially for (a) sarcasm classification in general; and (b) outer-circle varieties like en-IN for both the tasks, this benchmark will be useful to evaluate dialectal biases of language models, and contribute to research towards equitable performance across all varieties of English.
-Acknowledgments
-This work was funded by Google’s Research Scholar grant awarded in 2024.
+The overall difference in cross-domain performance is significantly higher for inner-circle varieties, compared to en-IN, indicating that fine-tuning with mixed-domain data may be less effective for such varieties.
+Across the varieties, however, en-UK, an inner-circle variety, performs better for google $\rightarrow$ reddit, whereas for the bert model, this is observed with the en-IN variety.
+Appendix E Detailed Feature Types
+We present detailed feature types identified in misclassified samples from different English varieties using the mistral model.
+Each table lists the feature, an example illustrating the feature, and the frequency count.
+Table 11 shows features found in en-AU samples, Table 12 shows features found in en-IN samples, and Table 13 shows features found in en-UK samples.
+Feature Example Count dial levelling of past tense verb forms when we order passionfruit cheecake and never get it 2 pronoun drop (We) also ordered 2 noodles, fish and dessert 3 there’s with plural subjects There’s been plenty of fossils and opals uncovered 1 no auxiliary in yes/no questions (Do) people still watch commercial TV?
+2 coll Hi Hospo workers afternoon when you wake up 28 cont Beige night 6
+Table 11: Features identified in the en-AU samples misclassified by mistral.
+Feature Example Count dial article omission Kamal is playing (a) cameo role 35 pronoun drop (I) can’t wait for animal park man 22 object/subject fronting Laws of India doesn’t apply on white man, he is above everything 12 ‘very’ as qualifier quantity is very very less 10 copula omission (I am) Waiting for cash to arrive in a tempo 18 coll It means benstokes 33 cont Alia Advani as her earlier name suits better.
+3 code Bridge chori ho jaaega 8
+Table 12: Features identified in the en-IN samples misclassified by mistral.
+Feature Example Count dial was for conditional were if it was in the autumn.
+1 group genitive their and their old boy network 1 levelling of past tense verb forms but could been more spicy.
+1 pronoun drop (It) took just 2 and a half minutes to lose his head 4 coll Warra calm and coherent interview.
+15 cont Haha yup!
+5
+Table 13: Features identified in the en-UK samples misclassified by mistral.